The weights of the Grade 11 students are known to be normally distributed with a mean of 54 kg and a standard deviation of 8 kg. Find the percentage of Grade 11 students with weights between 45 kg and 66 kg.



1
Expert's answer
2022-04-11T09:20:12-0400

We have a normal distribution, "\mu=54, \sigma=8."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}."


"z_1=\cfrac{45-54}{8}=-1.13;\\
z_2=\cfrac{66-54}{8}=1.50;\\
P(45<X<66)=P(-1.13<Z<1.50)=\\
=P(Z<1.50)-P(Z<-1.13)=\\
=0.9332-0.1292=0.8040 \text{ (from z-table).}"